Q.A projectile is fired on a horizontal ground.Coefficient of restitution between the projectile and the ground is e.Let a,b and c be the ratio of time of flight T(1)/T(2), max. height (H(1) / H(2) ) and horizontal range (R1/R2) in first two collisions with the ground.Then:(Options)


ย 


a. a = 1/e;


b= 1/e ;


c=1/e^2(e square)


ย or d. all of the above.

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Raj

Effect of cofficient of restitution can be calculated be vertical velocity say v effect.

Time of flight  proportional v

max height prportional to v^2

Haorizontal range prportional to v

T(1)/T(2)=1/e

H(1)/H(2)=1/e^2

R(1)/R(2)=1/e

So option a is correct

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..